Grade 12Mechanics

A pendulum consists of a small wooden block of mass m and length of thread l is suspended from a rigid support. A bullet of mass m1 is fired with a speed v. The bullet emerges out of the block with speed v/2 and the block just completes the vertical circle . Then velocity of the bullet v is?

Tanmayee
7 Years agoGrade 12

1 Answer

sree lakshmi
7 Years ago
on applying conservation of momentum 
m1v = m1(v/2) +mv’ 
mv’ = m1(v/2) …................(1)
for pendulum the KE is coverted into PE 
PE = mg2l 
KE = P2/2m  \Rightarrow (mv’)2/2m =[m1(v/2)]2/2m
KE=PE  \Rightarrow [m1(v/2)]2/2m = mg2l 
 [m1(v/2)] =4m2gl
(v/2)=4m2gl/m12 
v= 16 m2gl/m12
v =4m/m1\left ( \sqrt{gl} \right )